OpenStudy (anonymous):

WILL FAN AND MEDAL FOR HELP! (2 questions)

OpenStudy (anonymous):

The volume of the sphere is \[V=\frac{4}{3}\pi*r^3\]\[36\pi=\frac{4}{3}\pi*r^3\]\[\boxed{r=3~cm}\]Now the area of hte sphere is\[A=4*\pi*r^2\]\[A=4*\pi*3^2\]the answer is:\[\boxed{A=36\pi~cm^2}\]

OpenStudy (anonymous):

Considering the jar a cube\[V_{old}=a^3\]the new dimension will have the sides = 10a\[V_{new}=(10a)^3=1000a^3=\boxed{1000V_{old}}\]So the correct answer will be 1,000 times

OpenStudy (anonymous):

Okay, could you help with the two more?

OpenStudy (anonymous):

Fast ;)

OpenStudy (anonymous):

3. Calculate the surface are of a regular triangular pyramid with the base edges of length of 12 and a slant height of 12. (Hint: Remember that the base of a regular triangular pyramid must be an equilateral triangle, not necessarily congruent to the sides of the pyramid.) 4. Approximate the volume of the regular triangular pyramid in problem 3 if the height is approximately 11.5. (Hint: Remember that the base of a regular triangular pyramid must be an equilateral triangle, not necessarily congruent to the sides of the pyramid.)

OpenStudy (anonymous):

Answer choices for 3. a. 72 b. 216 c. 278.35 d. 288 Answer choices for 4. a. 239 b. 276 c. 717 d. 828